# Benchmarking LLM-based Machine Translation on Cultural Awareness

### v2

@@ 5 Experimental Settings
Methods in Comparison:
To fully evaluate the efficacy of LLM translations for cultural nuances, we compare the different prompting strategies (§4) on a tuning-free LLMs, as well as an open-sourced NMT model and a commercial MT system.
  • Prompting LLMs: We examine the different prompting strategies on OpenAI’s ChatGPT(gpt-3.5-turbo-1106) which is fine-tuned from GPT-3.5 with instructions. We also include LLaMA2-7B for comparison.
@@
  • NLLB: We use the NLLB 1.3B Team et al. (2022) model, which is a state-of-the-art multilingual MT model.
  • Commercial MT: We use the Google Translate engine in our comparison.
@@
6 Results and Analysis
Our analysis includes three parts: 1) automatic evaluations between LLM-based translation and popular NMT systems (§6.1) on six language pair translations; 2) assessing LLM prompting strategies using ChatGPT and LLaMA for En-to-Zh translation (§6.2); and 3) fine-grained human evaluation on a subset of En-to-Zh pairs (§6.3).
6.1 Overall Automatic Evaluation
(a) Traditional MT Systems
(b) LLM-based MT Systems
Figure 3: CSI-Match on Different Language Pair Translations: The coordinate scale range is from 40 to 100.
Model Overall-U NT-U Overall-U NT-U
EN-ZH ZH-EN
NLLB 22.4 17.7 6.38 0
NLLB-A 23.1 - 11.1 -
NLLB-R 22.9 - 17.3 -
LLaMA2 16.3 22.2 18.1 11.1
LLaMA2-A 20.2 - 29.3 -
LLaMA2-R 19.6 - 26.8 -
ChatGPT 55.8 61.9 54.4 72.2
Google 56.6 58.7 39.9 22.2
Table 4: Win Rates (%) of Understandability evaluated by GPT-4 for En-Zh Translation Pair: Overall-U means understandability on the entire dataset, and NT-U means non-translation CSIs’ understandability
Replacement and Append Work for NLLB and LLM-MT in CSI-Match.
@@ 6.1 Overall Automatic Evaluation
Figure 3 shows the CSI-Match scores (§3) for eight methods across cultural-related parallel corpora in six language pairs, which are collected by our pipeline (§2). Figure 2(b) shows the CSI-Match scores of traditional MT systems including vanilla NLLB, NLLB-A, NLLB-R, and Google Translate, whereas Figure 2(a) shows scores of LLM-based MT systems including LLaMA2, LLaMA2-A, LLaMA2-R and ChatGPT. We find that straightforward strategies using dictionaries of CSIs, such as Replace (the green solid line) and Append (the orange solid line), are effective on metrics that rely on string-matching, such as CSI-match, as well as other semantic matching metrics detailed in Appendix §G. However, the appending strategy significantly benefits LLaMA more than NLLB. This suggests that LLaMA’s capacity of in-context learning and instruction-following enables the flexible integration of cultural knowledge at test time, a capability not presented in traditional NMT systems like NLLB. Additionally, we measure the CSI-Match of ChatGPT using basic instructions (BI), and Google Translate (the red solid line). We find that Google Translate’s performance across different languages is generally more consistent compared to ChatGPT. Specifically, Google Translate performs better in non-Latin languages such as Chinese, Hindi, Tamil, and Telugu, when compared to ChatGPT.
Understandability Gap on Non-translation CSIs
Table 4 presents the understandability win rates assessed by GPT-4 for various MT systems compared to the reference in English-Chinese translation. Appendix §F has the detailed evaluation prompt. The evaluation metrics include both the overall win rates across the entire dataset (Overall-U) and the win rates of samples containing complex source items (CSIs) with no translations (NT-U). The results indicate that traditional terminology translation methods, such as Appending and Replacing, effectively improve the overall understandability across the entire dataset. However, they still encounter challenges in improving the comprehensibility of translations that contain CSIs without a well-established reference. Notably, ChatGPT performs better than any other MT systems, which potentially suggests that the instruction-tuning of LLMs beyond the translation task may be beneficial for the model to generate explanation-based translations for these non-translation CSIs.
6.2 Prompting Strategy Evaluation
Shot Method CSI-Match Overall-U NT-U
Zero BI 67.2 55.8 61.9
CT 84.0 59.4 -
CE 66.6 55.5 58.7
SE 65.1 55.7 58.7
SR 69.2 58.5 47.6
Two CE 67.1 57.7 73.0
SE 67.7 58.7 58.7
SR 68.2 56.6 57.1
Table 5: Evaluation of different strategies by ChatGPT on English-Chinese translations.
As LLM-based MTs open up the opportunity to incorporate free-form external knowledge for improving the understandability of non-translation CSIs, we explore various prompting strategies by incorporating additional cultural information alongside dictionaries to improve translation. We compare different prompting strategies on ChatGPT, focusing on English-to-Chinese translation. We examine both zero-shot and 2-shot approaches to obtain results from ChatGPT. Especially for complex prompting strategies including CE, SE, and SR, we use 2-shot examples that are tailored for culturally-aware translation. Table 5 shows the results of different prompting methods.
Enternal Knowledge Prompting
Comparing the strategy of using external knowledge in prompts (i.e., CT and CE), we observe that LLMs can effectively leverage both direct translations and indirect descriptions. Specifically, CT greatly enhances the CSI-Match score for CSI translation. While 2-shot CE demonstrates limited improvements on CSI-Match, it improves the overall understandability of the entire dataset. More importantly, 2-shot CE significantly improves the non-translation CSIs’ understandability to 73.0. This suggests that CSI explanations offer significant assistance in translating CSIs, especially for the ones without well-known translations. In Table 6, we quantitatively show the output examples of “Polenta,” an Italian corn porridge. Interestingly, its Chinese translation on Wikidata is merely a transliteration. Under the CT strategy, ChatGPT directly copies this transliteration into the output, which may be considered correct yet not comprehensible for native speakers of the target language. In contrast, ChatGPT using CE integrates an explanation into the translation, describing the CSI as “corn porridge” in Chinese. This translation is much easier for readers to understand the nature of Polenta.
Internal Knowledge Prompting
Next, we compare prompting strategies to elicit LLMs’ internal knowledge (i.e., SE, SR). We find that 2-shot SE improves translation performance on CSI-Match and Overall-U but doesn’t improve the NT-U, which suggests that ChatGPT itself still lacks sufficient knowledge of non-translation CSIs which are typically very unique to the target-language culture. On the other hand, SR proves beneficial on ChatGPT for CSI-Match but does not improve the understandability of CSI translations, indicating that LLMs inherently possess internal knowledge of low-frequency CSIs but still can not capture cultural nuances through the SR prompting strategy examined in this study. In Table 6, 2-shot SR translates “Polenta” as “cornmeal,” similar to the baseline (BI). However, SE successfully translates it into “corn porridge.” This suggests that ChatGPT possesses background knowledge of Polenta but struggles to spontaneously elicit it to generate and select the most understandable translations.
Strategy Outputs
BI 玉米面 (cornmeal)来自意大利北部,遍布阿根廷。
CT 波伦塔 (transliteration) 起源于意大利北部,在阿根廷非常常见。
CE (2-shot) 玉米粥 (corn porridge) 源自意大利北部,是阿根廷非常常见的食品。
SE (2-shot) 玉米粥 (corn porridge) 起源于意大利北部,是在阿根廷非常常见的食物。
SR (2-shot) 玉米面 (cornmeal) 来自意大利北部,在阿根廷非常常见。
Source Polenta comes from Northern Italy and is very common throughout
Argentina
Reference 玉米粥 (corn porridge)来自意大利北部,在整个阿根廷很常见
Knowledge Translation: 波伦塔 (transliteration)
@@ 6.2 Prompting Strategy Evaluation
Explanation: Polenta is a dish of boiled cornmeal that was historically
made from other grains. The dish comes from Italy. It may be served
served as a hot porridge.
Table 6: Output Examples of Prompting Strategy (Top) and a source-reference sentence pair with cultural knowledge for En-Zh translation (Bottom)
@@ 6.3 Human Evaluation
GPT-4 and Human Evaluation Are Consistent.
Table 7 shows the Overall-U evaluated by both GPT-4 and human annotators. Specifically, we randomly select 200 samples from the English-to-Chinese translation dataset and have human annotators evaluate the understandability of output translations by the four different methods. The results indicate a high level of consistency between human annotations and GPT-4’s evaluations
Method GPT-4 Overall-U Human Overall-U
NLLB 19.0 16.0
LLaMA2 13.0 14.0
ChatGPT 52.5 51.0
Google 54.0 55.5
Table 7: GPT-4 v.s. Human in Overall-U evaluation
Translation Strategies
To explore potential factors benefiting understandability, we let a human annotator examine the models’ translation strategies. In addition to the four strategies (mentioned in §3; Table 2), we propose three extra considerations 1) Copy that directly copies the source language of CSI into the target language; 2) Wrong that indicates entirely incorrect translations; and 3) Other that employs other strategies in translation. Figure 4 shows the ratio of each strategy in four MT systems. We find that models with higher understandability (i.e., ChatGPT and Google translate) use description and substitution at a significantly higher rate, indicating that these two strategies help improve the understanding of CSI for target-language speakers. Notably, LLaMA2 incorporates a higher frequency of substitution and description methods compared to traditional NLLB. However, this increased diversity in translation output comes at the cost of reduced stability in the outputs. As a result, LLaMA2 tends to yield more inaccurate translations, whereas NLLB relies more on Literal Translation and Transliteration to translate CSIs.
To compare the consistency between automatic and human evaluation, we also calculate the correlation between automatic evaluation metrics and human evaluation, where we consider the translation correct in human evaluation if the translation is categorized into all the translation strategies except Copy and Wrong. The Pearson’s correlation between the human evaluation and BLEU, BLEURT and COMET are 88.8, 90.5, and 89.5 respectively, whereas CSI-Match has the highest correlation score at 94.7 with human evaluation, suggesting that CSI-Match can be an efficient evaluation metric for CSIs with a reference translation.
Figure 4: Percentage of Translation Strategies
7 Related Works
Cultural-aware Machine Translation:
As languages and cultures are highly intertwined, there is a growing desire towards empowering cultural awareness of machine translation systems Nitta (1986); Ostler (1999); Hershcovich et al. (2022). However, as cultural nuances are subtle, collecting culturally sensitive data Akinade et al. (2023) remains costly and time-consuming. Besides, it is also challenging to perform a human-centered evaluation of the cultural nuances Liebling et al. (2022). Existing studies have proposed strategies to evaluate cultural awareness of traditional MT systems by grounding images Khani et al. (2021), adapting entities Peskov et al. (2021) or targeting at dilates Riley et al. (2022). Different from evaluating traditional MT systems, we focus on evaluating the cultural awareness of LLM-based translation.
MT with Terminology Dictionaries
Previous studies have proposed to integrate terminology dictionaries into NMT models Dinu et al. (2019) and LLM-based MT systems Ghazvininejad et al. (2023), proving effective for terminology translation. A major thrust of methods either modify the model architectures to integrate the dictionary or feed the translation dictionary as parts of inputs to NMT models. In this study, we focus on evaluating LLM-based MT methods with conventional NMT methods without modifying the underlining models, as the parameters of LLMs (e.g., ChatGPT) may not be accessible. Moreover, compared to traditional terminology translation from popular domains (e.g., finance, medicine), translating cultural-specific items carries its unique challenge. CSIs are unique to specific cultural groups, leading to difficulty in understanding the cultural nuance of CSIs for individuals from other cultures. Thus, beyond accuracy-focused metrics, assessing CSI understandability is also crucial and underexplored in the literature.
LLM-based MT:
Large language models, such as GPT-3 Brown et al. (2020), have proven effective in machine translation for various high-resource languages Hendy et al. (2023); Jiao et al. (2023). Particularly, a few recent studies have investigated the performance of LLM-based MT, including formality control of translation outputs Garcia and Firat (2022), in-context translation ability during pre-training Shin et al. (2022), and multilingual translation Scao et al. (2022). However, the exploration of LLM-based MT on culturally sensitive texts is still missing.
External Knowledge for MT:
There have been multiple threads of research efforts on integrating external knowledge such as bilingual translation lexicons for neural machine translation systems, including probability interpolation of lexicons Arthur et al. (2016); Khandelwal et al. (2021), data augmentation by back-translations Hu et al. (2019), decoding with a phrase memory Wang et al. (2017), and pre-training with an entity-based denoising objective Hu et al. (2022). Despite the effectiveness, these methods require further fine-tuning of the original MT systems, while we focus on tuning-free methods for integrating external knowledge for LLM-based MT in this study.
@@ 8 Conclusion
In this paper, we introduce a novel data curation pipeline aimed at constructing a culturally sensitive parallel corpus to assess the cultural awareness of MT systems. Additionally, we propose a reference-free metric to evaluate the understandability of translations for cultural-specific content by GPT-4. We also devise simple yet effective prompting strategies to enhance the understandability of LLM-based translations. Despite the effectiveness, several challenging questions remain open. First, it is non-trivial to incorporate cultural-specific information beyond a single entity such as discourse information. Besides, our prompting strategies leverage mostly external cultural knowledge in the form of texts. How to leverage multimodal knowledge from images and structured knowledge graphs to resolve cultural ambiguity deserves further investigation.
@@
Our work provides one step further toward understanding and evaluating the cultural awareness of LLM-based machine translation. We provide a culturally sensitive parallel corpus with rich annotations on cultural-specific items. Thereby, prompting strategies in our study, due to the cost ChatGPT API, is limited to the language pair (i.e., English-Chinese). To manually verify the quality of the data, we control the data size in a small one. We will provide our code repository to facilitate the future adaptation of our pipeline for more language pairs in diverse language families. Besides, as we focus on the evaluation of cultural-specific items in this study, the evaluation of cultural awareness beyond a single entity deserves further investigation. In addition to cultural-specific items, there are many other types of cultural errors that exist in the translation process, such as linguistic style and slang Hershcovich et al. (2022). Our work aims to mitigate cultural errors by starting from CSI, in order to promote advancements in cultural-aware machine translation datasets, models, and evaluation methods. This is of great importance for enabling machine translation to play a larger role in cross-cultural contexts. Further, we only try several promoting strategies in our study, which is because our work focuses on benchmarking the cultural awareness of current LLM-based MT systems, in the future, we’ll test other methods such as instruction tuning to improve the performance.
@@
Although our study designs a suite of simple but effective prompting strategies to enhance the cultural awareness of LLM-based machine translation, we still observe the weakness of LLM-based translation on cultural concepts in certain regions (e.g., Asia) and hallucinations on low-frequency entities. Potential usage of these LLM translation outputs may still result in misinformation spread. Before deploying our methods to create reliable content such as creating translations of Wikipedia articles, practitioners should ensure another round of human post-editing. In the future, we’ll make our dataset public to assist further research on cultural-awareness of MT systems. During the annotation process, the annotators (native speakers of the target languages) consist of the authors of this article, who know the goals of the study.

### v3

@@ 5 Experimental Settings
To fully evaluate the efficacy of MT system translating CSIs, we compare NMT systems with LLM-based MTs (§5.1). Secondly, to investigate the performances of traditional terminology translation methods on CSI translations, we evaluate two dictionary-based terminology methods on open-sourced NMT and LLMs (§5.2). Moreover, to gauge the capacity of LLMs to leverage external knowledge, we evaluate 4 cultural knowledge prompting strategies on tuning-free LLMs (§5.3).
5.1 MT systems in Comparison
We evaluate the following MT systems:
@@ 5.1 MT systems in Comparison
  • NMTs: We asses the NLLB 1.3B NLLB Team et al. (2022) model, which is a state-of-the-art multilingual MT model. We also use the Google Translate engine in our comparison.
  • LLMs: We examine ChatGPT (GPT-3.5-turbo-1106) and the open-source LLaMA2-7B for comparison, as both have been proven to be efficient multilingual MT tools Zhu et al. (2023).
5.2 Dictionary-based Methods in Comparison
For the open-sourced models (i.e., LLaMA and NLLB) we experiment with two additional methods proven to be highly effective in terminology MT during inference Dinu et al. (2019). Specifically, we employ 1) the Append method: append the CSI dictionary before the input, whose format is “<CSI1>:<CSI1 Translations>,…,<CSIn>:<CSIn Translations>[Source language]”; and 2) the Replace method: replace the CSIs in the source sentence with their translation in target language. For LLaMA2, we use the following prompt in 8 shots: [Source language]:[Source sentence] = [Target language]:[Target sentence], which is efficient for machine translation Zhu et al. (2023).
5.3 Prompting Strategies in Comparison
We explore various prompting strategies to introduce cultural knowledge into LLM-based MT. Our strategies generate in-context examples to integrate additional cultural knowledge, which involves employing CSI dictionaries and CSI explanations. Table 3 shows examples of four prompting strategies.
  • Basic Instruction (BI) The basic machine translation prompt of LLMs (e.g. ChatGPT).
  • External CSI Translation (CT) To assess the impact of incorporating a CSI dictionary within the prompts, we include CSIs along with their corresponding translations prior to the basic translation instruction BI.
  • External CSI Explanation (CE) CSIs may not have a direct equivalence in the target language’s culture. Therefore, it becomes necessary to translate based on the explanation of the CSI to assist the target audience in better understanding the content. To assess the impact of explanations, we include the CSI explanation obtained from Wikipedia in the prompt before the basic translation instructions.
  • Self-Explanation (SE) We also examine LLMs’ internal knowledge for explaining the meaning of CSIs. Inspired by Chain-of-Thought (CoT) Wei et al. (2022); Kojima et al. (2022), we treat the explanation of CSIs in a source sentence as an intermediate reasoning step before translating the whole sentence. We design the explanation prompting strategy in two steps for machine translation. First, we prompt the LLM to explain the meaning of all CSIs in the source sentence. Second, we ask the LLM to translate the whole sentence by combining the LLM’s explanation with another prompt instruction.
6 Results and Analysis
In this section, we 1) compare the CSI translation performances of LLM-based MT systems with that of NMT systems (§6.1); 2) evaluate dictionary-based terminology translation methods to explore if they work on CSI translations (§6.2); 3) compare four prompting strategies of LLM-based MTs to explore how different prompts affect the LLMs’ cultural awareness (§6.3); 4) conduct human evaluation to verify the correlation between automatic evaluation metrics and human assessment (§6.4).
6.1 Evaluating LLM-based MT v.s. NMT
To compare the cultural awareness of LLM-based MT systems with that of NMT systems, We employ two automatic metrics (CSI-Match and PTA) to evaluate 4 MT systems: the vanilla NLLB and Google Translate, and the BI prompting of LLaMA2 and GPT-3.5.
Figure 3: CSI-Match results on six language pairs.
NMTs Excel in CSI-Match on Low-resource Languages
We evaluate the accuracy of CSI translations using CSI-Match across six language pairs, as shown in Figure 3. For the two Romance languages (French and Spanish), the performances of four MT systems are quite similar. However, Google Translate generally exhibits more consistent performance in non-Romance languages compared to LLM-based MT. NLLB’s poor performance on EN-ZH is due to its limited translation capacity on EN-ZH, as validated in previous benchmarking work Aharoni et al. (2019); Akinade et al. (2023). Specifically, NLLB and Google Translate outperform LLaMA2 and GPT-3.5 in translating three Indian languages: Hindi, Tamil, and Telugu. Additionally, for low-resource languages like Tamil and Telugu, the translation performance of LLM-MTs remains limited on traditional translation metrics (see Table 10). Therefore, LLM-MTs cannot yet be considered efficient multilingual translation tools for these low-resource languages.
GPT-3.5 Produces Better Pragmatic Translation on CSIs with No Established Translations.
Given the cost of evaluation by commercial tools and human experts across all language pairs, we focus on English-Chinese in both translation directions when evaluating the pragmatic translation quality by PTA. Figure 4 presents the PTA assessed by GPT-4o for four MT systems’ output compared to the reference sentences. In addition to PTA across the entire dataset, we separately evaluate PTA of samples containing CSIs with no established translations (NT-PTA). Notably, GPT-3.5 performs better than any other MT systems on NT-PTA, which potentially suggests that the instruction-tuning of LLMs beyond the translation task may be beneficial for the model to generate free translations that are easily understood by the target culture, especially for non-translation CSIs.
Figure 4: PTA results on English-Chinese translations.
6.2 Evaluating Dictionary-based Methods
Model Method CSI-Match PTA
NLLB EN-ZH Vanilla 53.1 17.1
@@ 6.2 Evaluating Dictionary-based Methods
[…]
ZH-EN Vanilla 64.9 6.7
[…]
LLAMA2 EN-ZH Vanilla 45.0 12.1
[…]
ZH-EN Vanilla 70.9 12.0
[…]
LLaMA is More Robust at Leveraging CSI Dictionaries than NLLB.
@@
We evaluate two dictionary-based terminology methods on NLLB and LLaMA for English-Chinese translations, as shown in Table 4. We find that straightforward strategies using dictionaries of CSIs, such as Replace and Append are effective for both NLLB and LLaMA on metrics that rely on string-matching (i.e. CSI-Match), as well as other semantic matching metrics (see Table 10). However, the appending strategy significantly benefits LLaMA more than NLLB. This suggests that LLaMA’s ability for in-context learning and instruction-following enables the flexible integration of cultural knowledge at test time, a capability not present in traditional NMT systems like NLLB. Furthermore, traditional terminology translation methods can improve the PTA across the entire dataset. Without dictionaries, they still encounter challenges in improving the comprehensibility of translations that contain CSIs.
6.3 Evaluating Prompting Strategies
Model Method CSI-Match PTA NT-PTA
GPT-3.5 BI 66.2 33.2 31.7
[…]
LLaMA2 BI 43.7 11.3 11.1
[…]
Strategy Outputs PTA of GPT-4o PTA of Human
BI {CJK}UTF8gbsn就像意大利的polenta concia一样,它可以作为主菜食用。 Lose Lose
CT {CJK}UTF8gbsn和在意大利一样,波伦塔(transliteration) 在意大利被当作主菜。 Lose Lose
CE {CJK}UTF8gbsn就像意大利的奶酪玉米粥(cheese corn porridge)一样,它可以作为主菜食用。 Win Win
SE {CJK}UTF8gbsn就像意大利的奶酪玉米粥(cheese corn porridge)一样,它可以作为主菜食用。 Win Win
Explanation by GPT-3.5: Polenta is a traditional Italian dish that originated in Northern Italy. It is a
type of porridge made from cornmeal, and is similar in consistency to grits or cornmeal mush.
Source Just like polenta concia in Italy, it is eaten as a main dish.
Reference {CJK}UTF8gbsn就像意大利的玉米粥 (corn porridge)一样,它可以作为主菜食用。
Knowledge Translation: {CJK}UTF8gbsn波伦塔 (transliteration)
@@ 6.3 Evaluating Prompting Strategies
Explanation: Polenta is a dish of boiled cornmeal that was historically made from other grains.
The dish comes from Italy. It may be served served as a hot porridge.
Table 6: The output example of four prompting strategies on GPT-3.5 for En-Zh translation.
LLM-based MTs open up the opportunity to incorporate free-form external knowledge to enhance the pragmatic translation quality of CSIs, especially for those without dictionaries. We explore various prompting strategies by integrating additional cultural knowledge, including dictionaries and explanations, to improve translations. We compare different prompting strategies for English-to-Chinese translations, employing 2-shot prompting approaches to obtain results from GPT-3.5 and LLaMA2. Table 5 shows the evaluation results.
External CSI Knowledge Improves LLM-MT.
When comparing the strategies of using external knowledge in prompts (i.e., CT and CE), we observe that LLMs can effectively leverage both direct translations and indirect explanations. Specifically, CT enhances the CSI-Match score for CSI translations. However, CT is not effective when CSIs have no existing translations. In contrast, the 2-shot CE approach using GPT-3.5 improves NT-PTA from 31.7 to 41.3 in English-to-Chinese translations. This suggests that CSI explanations can notably aid in translating CSIs, particularly those without well-known translations. For LLaMA, the PTA score is similar to the baseline (with differences of fewer than 10 examples out of 778 data points) due to the limitations of LLaMA2-7B’s capacity for English-to-Chinese translations. However, CE and SE approaches with LLaMA still show an improvement in NT-PTA, indicating that LLaMA2-7B can also leverage external explanations to improve the translation quality of CSIs without existing direct translations.
LLMs’ CSI Explanations Also Help.
We use SE to elicit LLMs’ internal knowledge and find that the 2-shot SE approach with GPT-3.5 improves translation performance across all metrics for English-to-Chinese translations. This suggests that GPT-3.5 may already possess a significant amount of cultural knowledge about CSIs and can integrate this knowledge into the translation process. For LLaMA2, the PTA of SE is close to the baseline, and the improvement in NT-PTA is also limited, indicating that LLaMA2-7B may not have sufficient cultural knowledge of CSIs for English-to-Chinese translations.
Prompting with CSI Explanations Encourages LLMs to Do Free Translations.
In Table 6, we provide examples of the CSI “polenta”, an Italian corn porridge. Its Chinese translation on Wikidata is merely a transliteration. Under the CT strategy, GPT-3.5 directly copies this transliteration into the output, which may be considered correct but not comprehensible for native speakers of Chinese. In contrast, using the CE strategy, GPT-3.5 integrates the CSI explanation into the translation, freely translating the term “corn porridge” into Chinese. This makes it easier for readers to understand the nature of “polenta”. Furthermore, The SE strategy successfully generates an explanation for “polenta” and incorporates it into the translation as “corn porridge”, which leads to better comprehension for Chinese native speakers, as is reflected in the ratings of GPT-4o and the human annotator.
@@ 6.4 Human Evaluation
Metrics Human Acc. Human PTA
BLEU 79.6 86.2
BLEURT 80.6 86.6
COMET 77.8 89.3
Exact-Match 77.0 81.7
CSI-Match 88.7 90.0
GPT-4o PTA 87.1 95.7
Table 7: Pearson’s Coefficients between automatic metrics and human evaluation on CSI translation
To compare the consistency between automatic metrics and human evaluation, we conduct a human evaluation on a subset of 200 English-to-Chinese translations. We assess the outputs from eight MT systems: NLLB, LLaMA2, Google Translate, and GPT-3.5 across zero-shot BI, and two-shot BI, CT, CE, and SE settings. The native Chinese speaker evaluates the accuracy and PTA of the outputs. We then calculate Pearson’s correlation coefficients between automatic evaluation metrics and human assessments. Specifically, we compare the performance of CSI-Match and PTA with four traditional automatic evaluation metrics: BLEU, BLEURT, COMET, and Exact-Match. The results, presented in Table 7, indicate that CSI-Match exhibits the highest correlation with human accuracy, while GPT-4o PTA shows the highest correlation with human PTA. These findings suggest that CSI-Match and PTA are effective evaluation metrics for assessing the translation quality of CSIs, which can better capture cultural nuances than traditional metrics.
@@ 7 Conclusion
To advance culturally-aware machine translation, we curate a high-quality, diverse parallel corpus (CAMT) with rich CSI annotations in 6 language pairs using an automated pipeline. We introduce two evaluation metrics, CSI-Match and PTA, to assess translation quality concerning cultural nuances. Our evaluation of LLM-based MT and NMT systems using CAMT reveals that LLMs can effectively incorporate external cultural knowledge, enhancing the pragmatic translation quality of CSIs. Our work provides essential data sources and insights for advancing culturally-aware machine translation, laying the groundwork for future investigation in this field.
@@
Language Pairs in Evaluation
Our work takes a significant step toward toward understanding and evaluating the cultural awareness of machine translation on CSIs. We provide a culturally sensitive parallel corpus with rich annotations on cultural-specific items in six languages pairs. However, due to the cost of evaluation by commercial LLMs and human experts across all language pairs, we conduct parts of our experiments on English-Chinese translations, whose data quality is also verified by human experts. Building on our insights into English-Chinese translation, we hope to encourage future work to verify our findings on other language pairs, and we will release our code repository to streamline further investigations.
Cultural-Awareness Definition
In this study, we focus on the evaluation of cultural-specific items (CSIs). However, evaluating cultural awareness beyond individual entities also deserves further investigation. Besides CSIs, many other types of cultural errors persist in the translation process, such as those related to linguistic style and slang Hershcovich et al. (2022). Our work aims to mitigate cultural errors by starting with CSIs, promoting advancements in culturally-aware machine translation datasets, models, and evaluation methods. This is crucial for enabling machine translation to play a larger role in cross-cultural communications.
Evaluation by LLM
Recent research has shown that GPT-4 demonstrates a high correlation with human experts in evaluating generation performance Rafailov et al. (2023); Kocmi and Federmann (2023); Li et al. (2024). However, using GPT-4 as an evaluator may still pose fairness issues due to internal biases and unbalanced language capabilities of LLMs. In this study, we aim to advance beyond traditional semantic alignment evaluation metrics to assess pragmatic translation quality in English-Chinese translations using GPT-4. Further investigation is needed to improve GPT-4’s effectiveness as an translation evaluator.
Prompting strategies
We only try 4 prompting strategies in our study, due to our work’s focus on benchmarking the cultural awareness of current LLM-based MT systems. In the future, we’ll test other methods, such as instruction tuning, to improve the performance of LLM-based MT.
@@
Although our study designs a suite of simple but effective prompting strategies to enhance the cultural awareness of LLM-based machine translation, we still observe the weakness of LLM-based translation on cultural concepts in certain regions (e.g., Asia) and hallucinations on low-frequency entities. Potential usage of these LLM translation outputs may still result in the spread of misinformation. Before deploying our methods to create reliable content such as creating translations of Wikipedia articles, practitioners should ensure another round of human post-editing. During the annotation process, the annotators (native speakers of the target languages) consist of the authors of this article, who know the goals of the study clearly.
Acknowledgements
We sincerely appreciate the valuable feedback provided by our reviewers, which greatly helped to improve the manuscript. BY and JH are supported by the Wisconsin Alumni Research Foundation. MJ is partially supported by the National Science Foundation (IIS-2438420). The content is solely the responsibility of the authors and does not necessarily represent the official views of the National Science Foundation.